If B is added to C-3.75.3j, the result is a vector in the positive direction of the y axis, with a magnitude equal to that of C. , the result is a vector in What is the magnitude of 8 Units This answer has no units the tolerance is +/-2% Question Attempts: 1 of 12 used SAVE
Answer #1

If a vector is pointing along y direction, means it x component is zero. Also magnitude of B + C has to be equal to magnitude of C.

で (a+3.7) i + ( +5.3) Atko b+5.3 Sod be Same as tude o C ラ6+ 5. J(з.yt + (s of 3: (S.3 → b+S-3 = 6.463 (3.1) + ( 1-1637) = | 3, 8736